# Picocavity-enhanced Raman Spectroscopy of Physisorbed H2 and D2 Molecules

### v2

@@ Abstract
We report on tip-enhanced Raman scattering (TERS) of H2 and D2 molecules physisorbed within a plasmonic picocavity at 10 K. The intense Raman peaks resulting from the rotational and vibrational transitions are observed at sub-nanometer gap distances of the junction formed by an Ag tip and Ag(111) surface, where a picocavity-enhanced field plays a crucial role. A significant red-shift of the H–H stretch frequency is found as the gap distance decreases, while the D–D stretch frequency is unaffected. Density functional theory, path-integral molecular dynamics, and quantum anharmonic vibrational energy calculations suggest that this unexpected isotope effect is explained by a different molecular density between H2 and D2 on the surface.
@@
The adsorption of hydrogen molecules on solid surfaces is the first step in permeation for fuel storage [1], catalytic hydrogenation [2], hydrogen embrittlement [3], and nuclear-spin isomer conversion [4]. To understand and control the elementary processes in such physical and chemical phenomena, hydrogen adsorption has been intensively studied in surface science [5, 6, 7]. Specifically, hydrogen, the lightest molecule, exhibits unique quantum effects that differ from those of other molecules when it is physisorbed on a surface through van der Waals (vdW) interactions [8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. However, only a few methods are available for characterizing such physisorption systems as the weak adsorption energy ( meV) necessitates measurements at cryogenic temperatures. Low-temperature scanning tunneling microscopy (LT-STM) has been employed as a local characterization technique to investigate the dynamics of hydrogen molecules weakly adsorbed on surfaces [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. Several studies have observed characteristic features in conductance spectra, attributed to the rotational transition of molecular hydrogen in the STM junction [21, 22, 23, 24]. However, the interpretation remains controversial; the features within the corresponding energy range [30–60 meV (240–480 cm-1) for H2] could be attributed to configuration switching [18, 29, 30], to a molecule–substrate stretching mode [31], or to multiple phonon excitation [32]. Additionally, the stretching mode has never been observed in these studies.
@@
Raman spectroscopy is a promising technique for characterizing hydrogen molecules on surfaces, as it allows the observation of both rotational and vibrational transitions of homonuclear diatomic molecules. Although the detection demands exceptional sensitivity due to the intrinsically small cross-section, single-molecule detection can be achieved by tip-enhanced and surface-enhanced Raman spectroscopy (TERS/SERS) [33, 34, 35, 36, 37]. Recent studies have established that picocavities, atomic-scale protrusions present in plasmonic nanojunctions, play a crucial role in achieving extreme confinement of the electromagnetic field to Ångström scales [38, 39, 37], resulting in a significant enhancement of Raman scattering [40, 41, 42] and ultrahigh spatial resolution [43, 44, 45, 46]. Yet, the exceptional sensitivity in TERS/SERS has been demonstrated primarily for organic molecules chemisorbed onto the substrate, where the enhancement benefits from chemical effects arising from the molecule–surface (and molecule–tip) interactions [47]. Therefore, applying TERS/SERS to physisorbed systems has remained challenging.
@@
Here we demonstrate that LT-TERS can detect hydrogen molecules physisorbed onto an Ag(111) surface with an Ag tip [Fig. 1(a), inset], allowing the identification of their rotational and vibrational states. Furthermore, precise gap-distance control enables us to investigate not only the contribution of the picocavity field to the Raman scattering process but also the influence of tip proximity on the rotational/vibrational energies of the molecules through non-trivial nuclear quantum effects. Density functional theory (DFT), path-integral molecular dynamics (PIMD), and quantum vibrational energy calculations suggest that the tip–molecule–surface potential energy profiles are critically affected by the intermolecular and tip–molecular interactions, leading to unusual isotope effects on vibrational frequencies.
@@
The LT-TERS experiments were performed in an ultra-high vacuum chamber [see Sect. I-A of the Supplemental Material (SM) [48]]. The 532-nm incident laser beam was linearly polarized along the tip axis and focused on the apex of a focused-ion-beam-sharpened Ag tip [40, 42] using an in-situ parabolic mirror, and the scattering light was detected by an ex-situ spectrometer. We conducted DFT calculations [49, 50] of H2/D2 on a Ag(111) (periodic) surface, employing the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional and screened pairwise vdW interactions [51] (see SM Sect. I-B [48]). The preferred adsorption sites (fcc/hcp hollow) and binding energies (30 meV) were in very good agreement with previous studies of isolated molecules on the surface [52, 53]. Unless specified otherwise, we considered a coverage of 0.69 monolayer (ML), motivated by the results of PIMD simulations at LT (see SM Sect. I-B [48]). At 0.69 ML coverage, the intermolecular attractive interaction increases the binding energy by 55 meV compared to the lowest coverage of 0.06 ML considered here (Fig. S4). Over the sample, we placed the Ag tip model shown in SM Sect. I-B [48] (see also Fig. S5), which was also previously discussed in Refs. [46, 54, 41].
@@
H2 or D2 gas was introduced into the chamber where a clean Ag(111) surface was cooled at 10 K (see SM Sect. I-A [48]). Regardless of laser illumination, no static hydrogen molecule on the surface was imaged with LT-STM (Fig. S1). In contrast, clear peaks were observed in TERS spectra under these conditions. Figure 1(a) displays typical TERS spectra of H2 and D2 on Ag(111). For H2, intense peaks were observed at 351 and 4121 cm-1, along with a weak peak at 4443 cm-1. Similarly, for D2, intense peaks appeared at 169 and 2967 cm-1, with a weak peak at 3133 cm-1. The three peaks can be assigned to rotational/vibrotational modes of hydrogen. Note that broad or asymmetric peak shapes are artifacts in the wide wavenumber range spectra (see SM Sect. I-A [48]). The feature around 100 cm-1 common to both spectra originates from phonon modes of the Ag tip [41, 55], which is present even in the absence of hydrogen adsorption (Fig. S2). Figure 1(b) depicts the energy diagram of rotational/vibrational levels and allowed transitions under the Raman selection rule. The observed peaks, in order from lowest to highest, correspond to the lowest-energy rotational transition [; “rot” labeled in Fig. 1(a)], the lowest-energy vibrational transition [; “vib”], and their combination [; “rot+vib”], respectively. The three modes were also observed in high-resolution electron energy loss spectroscopy (HREELS) of hydrogen on a polycrystalline Ag film [8] and a Cu(001) surface [9] at approximately 10–12 K; however, better energy resolution (10 cm-1) in TERS than that of the HREELS (28 cm-1 [9]) makes the red-shifts from gas-phase hydrogen more discernible (Table S4). We should also note that only the stable nuclear spin isomer of each isotope, i.e., para-H2 and ortho-D2 [Fig. 1(b)], is detected in TERS (Fig. S2). On the surface, the stable isomer is dominant due to the ortho–para conversion [8, 4, 52] (see SM Sect. II-A [48]).
@@
The rotational transition at 351 cm-1 for H2/Ag(111) [Fig. 1(a)] is slightly red-shifted from 354 cm-1 in the gas phase [56]. We calculate the H–H bonding distance to be 75.4 pm on Ag(111) (cf. 75.1 pm in the gas phase). Considering a rigid-rotor approximation, the rotational transition energy of para-H2 is calculated to be 353 cm-1 on the surface, compared to 356 cm-1 in the gas phase. This is in good agreement with the experiment. Notably, the H–H bonding distance of H2 on Ag(111) varied by only 0.06 pm at the closest tip position with respect to simulations without any tip. This suggests that the hydrogen molecules on the surface and in the STM junction behave as nearly free rotors.
@@
In the TERS spectrum [Fig. 1(a)], the vibrational frequency of 4121 cm-1 exhibits a red-shifted relative to the gas-phase value of 4161 cm-1 [56]. The vibrational transition energy is calculated to be 4088 cm-1 on the surface (without the Ag tip), after scaling the calculations performed in the harmonic approximation (see SM Sect. I-C [48]; see also Fig. S9 for a simulated LT-TERS spectrum [46]). The slight discrepancy between the experimental and calculated frequencies are consistent with the typical spurious softening of the PBE functional. Compared to the calculated gas-phase value, the vibrational frequency shows a red-shift of 73 cm-1 upon the adsorption, which is larger than, but in reasonable agreement with the experiment.
@@
The TERS intensity is strongly correlated with the localized surface plasmonic resonance (LSPR) of the tip–sample junction. In general, two types of enhancement mechanisms contribute to TERS [36, 47]: electromagnetic enhancement through LSPR and chemical enhancement originating from charge transfer between the target molecule and the tip/surface. In physisorption systems, the electromagnetic enhancement is expected to be dominant.
@@
To elucidate the role of electromagnetic enhancement, it is crucial to investigate the gap-distance dependence of TERS signals. Figure 2(a) shows the waterfall plot of the gap-distance dependent TERS spectra recorded for D2/Ag(111). When the tip approaches the surface, the TERS intensities of the rotational and vibrational modes increase apparently exponentially [Fig. 2(b)], in agreement with the prediction based on the electric field of a light-induced dipole in the plasmonic tip [57]. A broad background with a maximum at 590 nm is attributed to electronic Raman scattering which represents the spectral response of the LSPR in the junction [58]. The gap-distance dependent intensities of both rotational and vibrational modes closely follow the plasmonic feature [Fig. 2(b)], indicating the dominant contribution of the gap-mode plasmon to the TERS enhancement. Additionally, we observed that changes in both the plasmon background and the molecular Raman peak occurred by an accidental modification of the tip-apex structure (Fig. S10), implying that the TERS enhancement is susceptible to the atomic-scale structure of the tip apex, which generates a plasmonic picocavity.
@@
The gap-distance dependence of the TERS intensities is well reproduced by simulating the electronic field distribution in the Ag-tip–Ag-surface junction using the finite element method (FEM) (see SM [48] for details). To model the plasmonic picocavity, we assume an atomic-scale protrusion (0.5-nm radius of curvature) attached to a blunt tip body [40]. The black curve in Fig. 2(c) shows the field enhancement at the expected position of a D2 molecule located beneath the Ag tip, indicating a steep increase as the tip approaches. We confirmed the critical role of this sub-nanometric structure in TERS; as shown by the orange curve in Fig. 2(c), the tip without the atomic-scale protrusion exhibits a more moderate increase of the field enhancement as the tip approaches, which does not fit the experimental data [Fig. 2(b)]. The red curve in Fig. 2(c) shows the field enhancement at a possible position of the nearest neighboring molecule [Fig. 2(c), inset], indicating much lower enhancement. Therefore, TERS can achieve ‘nearly’ single-molecule sensitivity, even for physisorption systems. The contribution of LSPR to the TERS intensity was further confirmed by our excitation-wavelength dependent measurements (SM Sect. II-B [48]); a strong TERS peak was observed only when the scattering light wavelength matches with the resonance range of the LSPR (Fig. S3).
@@
TERS reveals the sensitivity of the rotational/vibrational states to the interactions between a target molecule and its local environment. Figures 3(a)–3(d) show the gap-distance dependence of high-resolution TERS spectra focusing on the rotational/vibrational modes of H2/D2. Notably, only the vibrational frequency of H2 exhibits a significant red-shift as the gap distance decreases [Fig. 3(c)]; H2 shows a peak shift of 33 cm-1 with a 200 pm displacement of the tip, whereas is merely cm-1 for D2. Additionally, the peak width of H2 vibrational mode also increases as the gap distance decreases (see SM Sect. II-C [48] for details). In contrast, the peak positions of the rotational modes for both H2 [Fig. 3(a)] and D2 [3(b)] are almost unchanged, consistent with our calculations showing that the interatomic distance is independent of the tip height, as described above.
@@
The isotope dependence of the red-shift of the vibrational band far exceeds the scale of typical isotope effects. To understand this notable behavior, we calculated the potential energy profile along the rigid vertical displacement of the center of mass of one molecule between the surface and the tip apex [Fig. 4(a)], while varying the tip–sample distance [see profiles (i)–(iv) at different in Fig. 4(b)]. At each position, we calculated the H2 and D2 harmonic intramolecular vibrational frequency (see SM Sect. II-C [48]). The calculations indicate a pronounced coupling of the stretch coordinate with the molecular translation along , evidenced by a steep variation of along this coordinate (Fig. S8). Furthermore, this variation changes depending on the tip–sample distance, leading us to define . As detailed in SM Sect. I-C [48], for each , we constructed a model 2D Schrödinger equation in the coupled coordinates and , parameterized by and . The solution of this model gives us the expected intramolecular stretch vibrational transition energy for each isotope, fully accounting for the anharmonic coupling between these coordinates. We concluded that if both H2 and D2 probe the same well of [Fig. 4(b)], the isotope effect on can account at most for a factor 2 between H2 and D2 (Tables S2 and S3), in disagreement with the experiment (Fig. 3).
@@
We demonstrate that this discrepancy can be resolved by accounting for density differences between H2 and D2. At lower temperatures, the solids formed by H2 and D2 have different molar volumes due to nuclear quantum effects [zero-point energy (ZPE) swelling [59]; see also SM Sect. I-B [48]]. As a result, solid D2 is more than twice as dense as solid H2 at 4.2 K [60]. Because the interaction potential between the molecules and the surface exhibits almost no corrugation [52], we expect that H2 adsorbs on Ag(111) at 10 K more sparsely than D2. To explore a limiting low-density regime, we repeated the calculations discussed in the previous paragraph for a coverage of 0.06 ML of H2 [Fig. 4(c)]. As shown in Fig. 4(d), the reduced intermolecular interactions at the surface cause the minimum of along to lie closer to the tip instead of closer to the surface, as decreases [(i) to (iv)]. Consequently, the average position of the molecule center of mass [bullet points in Figs. 4(b) and (d)] also shifts towards the tip as decreases [dotted line in Fig. 4(d)]. Computing in this scenario with the 2D quantum model of and , we find a larger variation with than in the higher-density case (Table S2). As shown in Fig. 4(e), considering the higher-density scenario for D2 (the filled bullets) and the lower-density scenario for H2 (the empty bullets) can account for the large isotope effect in the red-shift of the stretch transition as the tip approaches. Figure 4(e) also shows the experimental values from the plots in Fig. 3 (the solid and dotted curves for H2 and D2, respectively; see also Fig. S11). Although the exact gap distance in STM experiment is hard to determine (see SM Sect. II-C [48]), the overlap of the experimental curves with the calculated plots in Fig. 4(e) shows good agreement. The remaining numerical discrepancies can be attributed to the limitation of the DFT level of theory (see SM Sect. I-B [48]), the exact tip apex geometry, and to the uncertainty in determining the exact local surface coverage. Nevertheless, our results highlight the crucial role of intermolecular interactions in accurately interpreting vibrational spectroscopy of hydrogen isotopes in nanojunctions, where nuclear quantum effects have a significant impact [13].
@@
In summary, with LT-TERS, we observed the rotational/vibrational transitions of local H2 and D2 molecules through the interaction with the picocavity field. Our results show that LT-TERS is applicable to weakly adsorbed molecules, offering deeper insights into the local structures, reactions, and dynamics of adsorbates, such as chemical reactions at active sites [61] and surface diffusion (spillover) in catalytic processes [62], at the single-molecule level.

### v3

@@ Abstract
We report on tip-enhanced Raman spectroscopy of H2 and D2 molecules physisorbed within a plasmonic picocavity at 10 K. The intense Raman peaks resulting from the rotational and vibrational transitions are observed at subnanometer gap distances of the junction formed by an Ag tip and an Ag(111) surface, where a picocavity-enhanced field plays a crucial role. A significant redshift of the H-H stretch frequency is observed as the gap distance decreases, while the D-D stretch frequency is unaffected. Density functional theory, path-integral molecular dynamics, and quantum anharmonic vibrational energy calculations suggest that this unexpected isotope effect is explained by a different molecular density between H2 and D2 on the surface.
@@
The adsorption of hydrogen molecules on solid surfaces is the first step in permeation for fuel storage [1], catalytic hydrogenation [2], hydrogen embrittlement [3], and nuclear-spin isomer conversion [4]. To understand and control the elementary processes in such physical and chemical phenomena, hydrogen adsorption has been intensively studied in surface science [5, 6, 7]. Specifically, hydrogen, the lightest molecule, exhibits unique quantum effects that differ from those of other molecules when it is physisorbed on a surface through van der Waals interactions [8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. However, only a few methods are available for characterizing such physisorption systems as the weak adsorption energy ( meV) necessitates measurements at cryogenic temperatures. Low-temperature scanning tunneling microscopy has been employed as a local characterization technique to investigate the dynamics of hydrogen molecules weakly adsorbed on surfaces [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. Several studies have observed characteristic features in conductance spectra, attributed to the rotational transition of molecular hydrogen in the scanning tunneling microscopy (STM) junction [21, 22, 23, 24]. However, the interpretation remains controversial; the features within the corresponding energy range [30–60 meV (240–480 cm-1) for H2] could be attributed to configuration switching [18, 29, 30], to a molecule–substrate stretching mode [31], or to multiple phonon excitation [32]. Additionally, the stretching mode has never been observed in these studies.
@@
Raman spectroscopy is a promising technique for characterizing hydrogen molecules on surfaces, as it allows the observation of both rotational and vibrational transitions of homonuclear diatomic molecules. Although the detection demands exceptional sensitivity due to the intrinsically small cross section, single-molecule detection can be achieved by tip-enhanced Raman spectroscopy (TERS) and surface-enhanced Raman spectroscopy (SERS) [33, 34, 35, 36, 37]. Recent studies have established that picocavities, atomic-scale protrusions present in plasmonic nanojunctions, play a crucial role in achieving extreme confinement of the electromagnetic field to angstrom scales [38, 39, 37], resulting in a significant enhancement of Raman scattering [40, 41, 42] and ultrahigh spatial resolution [43, 44, 45, 46]. Yet, the exceptional sensitivity in TERS and SERS has been demonstrated primarily for organic molecules chemisorbed onto the substrate, where the enhancement benefits from chemical effects arising from the molecule–surface (and molecule–tip) interactions [47]. Therefore, applying TERS and SERS to physisorbed systems has remained challenging.
@@
Here, we demonstrate that low-temperature TERS (LT-TERS) can detect hydrogen molecules physisorbed onto an Ag(111) surface with an Ag tip [Fig. 1(a), inset], allowing for the identification of their rotational and vibrational states. Furthermore, precise gap-distance control enables us to investigate not only the contribution of the picocavity field to the Raman scattering process but also the influence of tip proximity on the rotational and vibrational energies of the molecules through nontrivial nuclear quantum effects. Density functional theory (DFT), path-integral molecular dynamics, and quantum vibrational energy calculations suggest that the tip–molecule–surface potential energy profiles are critically affected by the intermolecular and tip–molecular interactions, leading to unusual isotope effects on vibrational frequencies.
@@
The LT-TERS experiments were performed in an ultrahigh vacuum chamber [see Sec. I-A of Supplemental Material (SM) [48]]. The 532-nm incident laser beam was linearly polarized along the tip axis and focused on the apex of a focused-ion-beam-sharpened Ag tip [40, 42] using an in situ parabolic mirror, and the scattering light was detected by an ex situ spectrometer. We conducted DFT calculations [50, 51] of H2/D2 on a Ag(111) (periodic) surface, employing the Perdew-Burke-Ernzerhof exchange-correlation functional and screened pairwise van der Waals interactions [52] (see SM Sec. I-B [48]). The preferred adsorption sites (fcc or hcp hollow) and binding energies (30 meV) were in very good agreement with previous studies of isolated molecules on the surface [55, 65]. Unless specified otherwise, we considered a coverage of 0.69 monolayer (ML), motivated by the results of path-integral molecular dynamics simulations at low temperature (see SM Sec. I-B [48]). At 0.69 ML coverage, the intermolecular attractive interaction increases the binding energy by 55 meV compared to the lowest coverage of 0.06 ML considered here (Fig. S3). Over the sample, we placed the Ag-tip model shown in SM Sec. I-B [48] (see also Fig. S4), which was also previously discussed in Refs. [46, 54, 41].
@@
H2 or D2 gas was introduced into the chamber where a clean Ag(111) surface was cooled at 10 K (see SM Sec. I-A [48]). Regardless of laser illumination, no static hydrogen molecules on the surface were imaged with STM (Fig. S1). In contrast, clear peaks were observed in TERS spectra under these conditions. Figure 1(a) displays typical TERS spectra of H2 and D2 on Ag(111). For H2, intense peaks were observed at 351 and 4121 cm-1, along with a weak peak at 4443 cm-1. Similarly, for D2, intense peaks appeared at 169 and 2967 cm-1, with a weak peak at 3133 cm-1. The three peaks can be assigned to rotational and vibrotational modes of hydrogen. Note that broad or asymmetric peak shapes are artifacts in the wide wavenumber range spectra (see SM Sec. I-A [48]). The feature around 100 cm-1 common to both spectra originates from phonon modes of the Ag tip [41, 66], which is present even in the absence of hydrogen adsorption (Fig. S2). Figure 1(b) depicts the energy diagram of rotational and vibrational levels and allowed transitions under the Raman selection rule. The observed peaks, in order from lowest to highest, correspond to the lowest-energy rotational transition [; “rot” labeled in Fig. 1(a)], the lowest-energy vibrational transition [; “vib”], and their combination [; “rot+vib”], respectively. The three modes were also observed in high-resolution electron energy loss spectroscopy of hydrogen on a polycrystalline Ag film [8] and a Cu(001) surface [9] at approximately 10–12 K; however, better energy resolution (10 cm-1) in TERS than that of the electron energy loss spectroscopy (28 cm-1 [9]) makes the redshifts from gas-phase hydrogen more discernible (Table S4). We should also note that only the stable nuclear-spin isomer of each isotope, i.e., para-H2 and ortho-D2 [Fig. 1(b)], is detected in TERS (Fig. S2). On the surface, the stable isomer is dominant due to the ortho–para conversion [8, 4, 55] (see SM Sec. II-A [48]).
@@
The rotational transition at 351 cm-1 for H2/Ag(111) [Fig. 1(a)] is slightly redshifted from 354 cm-1 in the gas phase [62]. We calculate the H-H bond length to be 75.4 pm on Ag(111) (cf. 75.1 pm in the gas phase). Considering a rigid-rotor approximation, the rotational transition energy of para-H2 is calculated to be 353 cm-1 on the surface, compared to 356 cm-1 in the gas phase, showing good agreement with the experiment. Notably, the H-H bonding distance of H2 on Ag(111) varied by only 0.06 pm at the closest tip position with respect to simulations without any tip. This suggests that the hydrogen molecules on the surface and in the STM junction behave as nearly free rotors.
@@
In the TERS spectrum [Fig. 1(a)], the vibrational frequency of 4121 cm-1 exhibits a redshift relative to the gas-phase value of 4161 cm-1 [62]. The vibrational transition energy is calculated to be 4088 cm-1 on the surface (without the Ag tip), after applying a scale factor to the calculations performed in the harmonic approximation (see SM Sec. I-C [48]; see also Fig. S8 for a simulated LT-TERS spectrum [46]). The slight discrepancy between the experimental and calculated frequencies is consistent with the typical spurious softening of the Perdew-Burke-Ernzerhof functional. Compared to the calculated gas-phase value, the vibrational frequency shows a redshift of 73 cm-1 upon the adsorption, which is larger than, but in reasonable agreement with the experiment.
@@
The TERS intensity is strongly correlated with the localized surface plasmonic resonance (LSPR) of the tip–sample junction. In general, two types of enhancement mechanisms contribute to TERS [36, 47]: electromagnetic enhancement through LSPR and chemical enhancement originating from charge transfer between the target molecule and the tip or surface. In physisorption systems, the electromagnetic enhancement is expected to be dominant.
@@
To elucidate the role of electromagnetic enhancement, it is crucial to investigate the gap-distance dependence of TERS signals. Figure 2(a) shows the waterfall plot of the gap-distance-dependent TERS spectra recorded for D2/Ag(111). When the tip approaches the surface, the TERS intensities of the rotational and vibrational modes increase apparently exponentially [Fig. 2(b)], in agreement with the previously reported analytical model simplifying a plasmonic tip as a point dipole [60]. A broad background with a maximum at 590 nm is attributed to electronic Raman scattering which reflects the spectral response of the LSPR in the junction [67]. The gap-distance-dependent intensities of both rotational and vibrational modes closely follow the plasmonic feature [Fig. 2(b)], indicating the dominant contribution of the gap-mode plasmon to the TERS enhancement. Additionally, we observed that changes in both the plasmon background and the molecular Raman peak occurred by an accidental modification of the tip-apex structure (Fig. S11), implying that the TERS enhancement is susceptible to the atomic-scale structure of the tip apex, which generates a plasmonic picocavity.
@@
The gap-distance dependence of the TERS intensities is well reproduced by simulating the electronic field distribution in the Ag-tip–Ag-surface junction using the finite element method (see SM Sec. I-E [48]). To model the plasmonic picocavity, we assume an atomic-scale protrusion (0.5-nm radius of curvature) attached to a blunt tip body (Fig. S9). The black curve in Fig. 2(c) shows the field enhancement at the expected position of a D2 molecule located beneath the Ag tip, indicating a steep increase as the tip approaches. We confirmed the critical role of this subnanometric structure in TERS; as shown by the orange curve in Fig. 2(c), the tip without the atomic-scale protrusion exhibits a more moderate increase of the field enhancement as the tip approaches, which does not reproduce the experimental data [Fig. 2(b)]. The red curve in Fig. 2(c) shows the field enhancement at a possible position of the nearest neighboring molecule [Fig. 2(c), inset], indicating much lower enhancement. Therefore, TERS can achieve “nearly” single-molecule sensitivity, even for physisorption systems. The contribution of LSPR to the TERS intensity was further confirmed by our excitation-wavelength dependent measurements (SM Sec. II-B [48]); a strong TERS peak was observed only when the scattering light wavelength matches with the resonance range of the LSPR (Fig. S10).
@@
TERS reveals the sensitivity of the rotational and vibrational states to the interactions between a target molecule and its local environment. Figures 3(a)–3(d) show the gap-distance dependence of high-resolution TERS spectra focusing on the rotational and vibrational modes of H2 and D2. Notably, only the vibrational frequency of H2 exhibits a significant redshift as the gap distance decreases [Fig. 3(c)]; H2 shows a peak shift of 33 cm-1 with a 200 pm displacement of the tip, whereas is merely cm-1 for D2. Additionally, the peak width of the H2 vibrational mode also increases as the gap distance decreases (see SM Sec. II-C [48] for details). In contrast, the peak positions of the rotational modes for both H2 [Fig. 3(a)] and D2 [3(b)] are almost unchanged, consistent with our calculations showing that the interatomic distance is independent of the tip height, as described above.
@@
The isotope dependence of the redshift of the vibrational band far exceeds the scale of typical isotope effects. To understand this anomalous behavior, we calculated the potential energy profile along the rigid vertical displacement of the center of mass of one molecule between the surface and the tip apex [Fig. 4(a)], while varying the tip–sample distance [see profiles (i)–(iv) at different in Fig. 4(b)]. At each position, we calculated the H2 and D2 harmonic intramolecular vibrational frequency (see SM Sec. I-C [48]). The calculations indicate a pronounced coupling of the stretch coordinate with the molecular translation along , evidenced by a steep variation of along this coordinate (Fig. S7). Furthermore, this variation changes depending on the tip–sample distance, leading us to define . As detailed in SM Sec. I-C [48], for each , we constructed a model 2D Schrödinger equation in the coupled coordinates and , parameterized by and . The solution of this model gives us the expected intramolecular stretch vibrational transition energy for each isotope, fully accounting for the anharmonic coupling between these coordinates. We concluded that if both H2 and D2 probe the same potential energy profile [Fig. 4(b)], for reasonable position-dependent stretch frequencies the calculated peak shifts are in a ratio of at most 2:1 for H2:D2 (Tables S2 and S3), in disagreement with the measured 16:1 ratio (Fig. 3).
@@
We demonstrate that this discrepancy can be resolved by accounting for density differences between H2 and D2. At lower temperatures, the solids formed by H2 and D2 have different molar volumes due to nuclear quantum effects [zero-point energy (ZPE) swelling [68]; see also SM Sec. I-B [48]]. As a result, solid D2 is more than twice as dense as solid H2 at 4.2 K [69]. Because the interaction potential between the molecules and the surface exhibits almost no corrugation [55], we expect that H2 adsorbs on Ag(111) at 10 K more sparsely than D2. To explore a limiting low-density regime, we repeated the calculations discussed in the previous paragraph for a coverage of 0.06 ML of H2 [Fig. 4(c)]. As shown in Fig. 4(d), the reduced intermolecular interactions at the surface cause the minimum of along to lie closer to the tip instead of closer to the surface, as decreases [(i) to (iv)]. Consequently, the average position of the molecule center of mass [bullet points in Figs. 4(b) and (d)] also shifts toward the tip as decreases [dotted line in Fig. 4(d)]. Computing in this scenario with the 2D quantum model of and , we find a larger variation with than in the higher-density case (Table S2). As shown in Fig. 4(e), considering the higher-density scenario for D2 (the filled bullets) and the lower-density scenario for H2 (the empty bullets) can account for the large isotope effect in the redshift of the stretch transition as the tip approaches. Figure 4(e) also shows the experimental values from the plots in Fig. 3 (the solid and dotted curves for H2 and D2, respectively; see also Fig. S12). Although the exact gap distance in STM experiment is hard to determine (see SM Sec. II-C [48]), the overlap of the experimental curves with the calculated plots in Fig. 4(e) shows good agreement. The remaining numerical discrepancies can be attributed to the limitation of the DFT level of theory (see SM Sec. I-B [48]), the exact tip-apex geometry, and to the uncertainty in determining the exact local surface coverage. Nevertheless, our results highlight the crucial role of intermolecular interactions in accurately interpreting vibrational spectroscopy of hydrogen isotopes in nanojunctions, where nuclear quantum effects have a significant impact [13].
@@
In summary, with LT-TERS, we observed the rotational and vibrational transitions of local H2 and D2 molecules through the interaction with the picocavity field. Our results demonstrate that LT-TERS is applicable to weakly adsorbed molecules, offering deeper insights into the local structures, reactions, and dynamics of adsorbates, such as chemical reactivity at active sites [70] and surface diffusion (spillover) in catalytic processes [71], at the single-molecule level.